Two lenses of power –15 D and +5 D are in contact with each other. The focal length of the combination is ?

Grade:12

1 Answers

Apoorva Arora IIT Roorkee
9 years ago
The focal length fo a combination is given by
\frac{1}{F}=\frac{1}{f_{1}}+\frac{1}{f_{2}}
where\frac{1}{F}=P
so
\frac{1}{F}=-15+5=10D
F=1/10=0.1 cm
Thanks and Regards
Apoorva Arora
IIT Roorkee
askIITians Faculty